how can i solve using substitution?
x=y+4
2x+3y=43

Answer by stanbon(75887)

how can i solve using substitution?
x=y+4
2x+3y=43
-----
Solve for "x" and solve for "y":
2(y+4) + 3y = 43
2y+8 + 3y = 43
5y = 35
y = 7
==========
solve for "x":
x = y+4
x = 7+4
x = 11
-------------
Cheers,
Stan H.